mass of a proton is 1.67*10^-24 gram. what is the mass of 1,000 protons?


1
Expert's answer
2021-02-01T03:53:00-0500

mass of a proton = 1.67×10-24

Mass of 1000 proton = 1.67×10-24×1000

= 1.67×10-21